A 12.0 µF capacitor is charged by a 13.0 V battery through a resistance R. The capacitor reaches a potential difference of 4.00 V at a time 3.00 s after charging begins. What is the value of R.
Answer #1
   Voltage across a capacitor C chanrgingthrough a resistance R is
   V   =   V0* ( 1 - e-t/RC)
   4.00   =   13.0 * (1 - e3.00 / RC)
   e-3.00 /RC   =   1 - 4.00 / 13.0
   e3.00 /RC   =   13.0 / 9.00
   taking log on both sides
   3.00 /RC   =   loge(1.444)
   R   =   3.00 /0.3677 * 12.0 * 10-6
         =   6.799* 105   Ω
          =   679.9   kΩ